The number of times a person blinks was studied for a sample of 35 people. The sample's average was 47 blinks per minute. The population standard deviation is 7.8 blinks per minute. What is the 99% confidence interval for the population average?
Answer by stanbon(75887)

The number of times a person blinks was studied for a sample of 35 people. The sample's average was 47 blinks per minute. The population standard deviation is 7.8 blinks per minute. What is the 99% confidence interval for the population average?
-----------------
xbar = 47
ME = inbT(0.995 with df=34)
= 2.7284*1.3184 = 3.5972
--------------------
99% CI: 47-3.5972 < u < 47+3.5972
==========================================
Cheers,
Stan H.
=============